a store sells small notebooks for $6 and large notebooks for $10. if a student buys 6 notebooks and spends $48 how many of each size did he buy?
Answer by josgarithmetic(39620)

Let x = how many small
Let y = how many large

SIX NOTEBOOKS: x+y=6
SPENT $48: 6x+10y=48


Various ways to solve. One method can be, simplify the money equation.
3x+5y=24.
Multiply the notebook count equation by 3 to get 3x+3y=18 and subtract this equation from the simplified money equation:
(3x+5y)-(3x+3y)=24-18
and solve .... first for y.
...!
Can you do that, and then continue for finding x.